HUBBLE SPACE TELESCOPE DAILY REPORT #832 PERIOD COVERED: 0800 EST Daily Status Report as of 057/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of PC 3603 (Snapshot of Protoplanetary Nebulae) The Planetary Camera was used to make observations of the protoplanetary system CPD 5D573. All the observations were complete in course track. One of the observations was reported to have no target visible in the field of view. 1.2 Completed FOS/RD 4495 (FOS GIMP Test) The Geomagnetically Induced Movement Problem (GIMP) was successfully completed. (See Operations Note 2.4) All of the observations were received and displayed. The first set of observations were wave length calibrations and they were reported to look good. The spectral lines lined up nicely throughout the time the observations were made. The TALRDs in the second set of observations also looked good, with the LED image in the same position throughout the observations. The final group of observations were wave length calibrations, and they also look good. 1.3 Completed FOC/96 3790 (Polarization Imaging of Radio Galaxies) The Faint Object Camera (f/96) was used with filter F320W to make three ultraviolet images of the radio galaxy 3C225B. All three observations were executed in fine lock. There was one recentering event which occurred on the first observation. There were no targets visible in any of the observations. 1.4 Completed GHRS/2 3614 ( boron Probe of Stellar Structure and Mass Loss) The High Resolution Spectrograph was used to make four spectral observations of the abundance of Boron in star HD23249. The baseline guide star acquisition was made in single star course track using FGS #3. The onboard target acquisition returned to brightest map in the sixth map. There was no peakup needed. The small science aperture was then move to the target and the spectral observations were started on time. No problems were reported. 1.5 Completed Two Sets of WFC 4713 (WF/PC Calibration: Internal; Delta Flats) The Wide Field Camera was used to make a series of internal flat observations. The observations were monitored in the OSS, and no problems were reported. 1.6 Completed WFC 3253 (Optical and UV Observations of Radio Pulsars) The Wide Field Camera was used to make two observations of the Vela radio pulsar (aka, PSR 0833-45). Both observations were taken in fine lock. The observations were received and displayed. No problems were reported. 1.7 Completed PC 4140 (WF/PC Calibration: Rapid Internal Monitor Part 2) The Planetary Camera was used to make one internal flat observation to check for new forms of contamination and the return of quantum efficiency hysteresis. The observations was made during LOS, and still remains on the science tape recorder. It will be analyzed later. 1.8 Completed Three Sets of PC 4714 ( WF/PC Calibration: PC Internal Delta Flats) and Initiate a Fourth The Planetary Camera was use d to make internal flat observations using filter F673N. This image will be used in the restoration of Earth flats, which are altered due to changes in quantum efficiency which occurred after decontamination operations. The OSS monitored the observations, and no problems were reported. 1.9 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.10 Completed HSP/POL 4209 (HSP Detector Sensitivity Test) The High Speed Photometer detectors were calibrated using observations from standard star VID998. The observation was executed as scheduled, and monitored by the OSS. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 4 Successful : 4 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 13 Successful : 12 (056/2209Z) 2.3 Significant HSTARs: None 2.4 Operations Note: A Geomagnetic Induced Movement Problem (GIMP) Test was successfully completed yesterday. The Red Side test was completed (056/0500Z-0930Z), and the Blue Side test at (056/11Z-1600Z). The test results were displayed in the OSS, and the image motion has been successfully removed. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Two Sets of PC 4714 (WF/PC Calibration: Internal Delta Flats) 3.2 Complete GHRS/2 3584 (Probing the Voids II) 3.3 Complete GHRS/2 4067 (GHRS 1ST Order Grating Spectral Lamp Test) 3.4 Complete Two Sets of WFC 4713 (WF/PC Calibration: Internal Delta Flats) 3.5 Complete HSP/POL 4209 (HSP Detector Sensitivity Test) 3.6 Complete HSP/UV2 1095 (Variability of High Luminosity Stars) 3.7 Complete FOC/96 3591 (Massive Stars in Starburst Galaxies- FOC Images Part 1) 3.8 Complete WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) 3.9 Complete Two Sets of FOS/RD 3744 (Spectrophotometry of Phobos and Deimos) 3.10 Complete Two Sets of PC 3603 (Snapshot of Protoplanetary Nebulae) 3.11 Complete PC 4347 (WF/PC Calibration: Flood Retention Test) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-